OpenStudy (cherry1214):

Use the quadratic formula to solve the equation. If necessary round to the nearest hundredth. x^2+3=-4x

OpenStudy (mjdennis):

OK, what's the quadratic equation? And do you have the problem in the correct form to use it?

OpenStudy (cherry1214):

I don't know what the quadratic equation is

OpenStudy (cherry1214):

Or if its in the correct form

OpenStudy (mjdennis):

Seems lie you are on a bad path, trying to answer problems when you have not learned the material. Quadratic equation is a foundation in math. Hang on.

OpenStudy (cherry1214):

Ok

OpenStudy (mjdennis):

Sorry, I meant the quadratic _formula_.

OpenStudy (cherry1214):

i know

OpenStudy (mjdennis):

Quadratic formula is a shortcut to finding answers to problems, when the _form_ of the problem is a*x^2 + b*x + c = 0.

OpenStudy (cherry1214):

So what do I do then?

OpenStudy (mjdennis):

Do you have enough algebra to put your problem into this format, and tell me what a, b, and c are, while I the next step ready?

OpenStudy (cherry1214):

a-0.02 b-2.3 c-6

OpenStudy (cherry1214):

its already in that format

OpenStudy (mjdennis):

Well, no it isn't in that form. There is a '-4x' on the right side of the equal sign, and in the format, there is supposed to be a '0' on the right side and all the x^2 and x terms on the left. and none of your a,b,c are correct. Is this the equation we are working on: x^2 + 3 = -4x ???

OpenStudy (cherry1214):

yeah i was wrong

OpenStudy (cherry1214):

oh

OpenStudy (cherry1214):

wait idk

OpenStudy (cherry1214):

u had the right equation I didn't lol

OpenStudy (mjdennis):

NP. how can we move the '4x' part over the left side?

OpenStudy (cherry1214):

subtract?

OpenStudy (cherry1214):

on both sides

OpenStudy (mjdennis):

Sure. Just be careful with signs, we are subtracting 'negative four x' from both sides, like this: x^2 + 3 = -4x x^2 + 3 - (-4x) = (-4x) - (-4x) Your turn. Finish up the math on this step and let me know.

OpenStudy (cherry1214):

x^2+3-(-4x)=0

OpenStudy (mjdennis):

Now, what can you do when you have two minus signs in a row, like in '-(-4x)' ?

OpenStudy (cherry1214):

make it positive

OpenStudy (mjdennis):

Then put the terms in the right order: X^2 first, then x , then constants (numbers without any x).

OpenStudy (mjdennis):

and wite it all out for us

OpenStudy (cherry1214):

x^2+4x+3=0

OpenStudy (cherry1214):

is that right?

OpenStudy (mjdennis):

Yep, so tell me what a, b, and c are, while I the next step ready.

OpenStudy (cherry1214):

a-x^2 b-4x c-3

OpenStudy (mjdennis):

Sorry to take so long between some answers, my connection is really crazy today.

OpenStudy (cherry1214):

Its ok

OpenStudy (mjdennis):

Close, but look again at the format: a*x^2 + b*x + c = 0 a does not include the x^2 part, and b does not include the x part.

OpenStudy (cherry1214):

a-1 b-4 c-3

OpenStudy (mjdennis):

Next step uses this formula to get the answers:\[x= \frac{ -b \pm \sqrt{b ^{2} - 4ac} }{ 2a }\] Just want to check, do you know what \[\sqrt{(something)}\] and \[\pm (something)\] mean?

OpenStudy (cherry1214):

top one is square root don't know the bottom

OpenStudy (mjdennis):

NP. Correct on square root, that is the tough one. The other, it looks like you stitched together a 'plus' sign and a 'minus' sign, and that is exactly what it means. It means the formula has to be used twice, once like this:\[x= \frac{ -b + \sqrt{b ^{2} - 4ac} }{ 2a }\] and again like this: \[x= \frac{ -b - \sqrt{b ^{2} - 4ac} }{ 2a }\]

OpenStudy (cherry1214):

ugh

OpenStudy (mjdennis):

I know. For algebra, quadratic formula is about as messy as it gets. Lots of work.

OpenStudy (mjdennis):

You go plug in those numbers, and I'll do my copy, and we'll check your answers.

OpenStudy (cherry1214):

\[-4x \pm \sqrt{4^2-4(1*3)}\]

OpenStudy (cherry1214):

over 2(1)

OpenStudy (mjdennis):

Did you leave the 'x' in by mistake? Otherwise, yes, correct.

OpenStudy (cherry1214):

yea

OpenStudy (cherry1214):

Now what?

OpenStudy (mjdennis):

NP, we are trying to speak three 'languages', between math formulas and English and keyboarding! Finish the math and I can check your work!

OpenStudy (cherry1214):

ok

OpenStudy (cherry1214):

-7.6

OpenStudy (mjdennis):

Because we have an x^2, term, there will _always_ be two possible answers. (with two exceptions. Sometimes the part under the square root sign is equal to zero, and so you really have the same answer twice. And sometimes the part under the square root is negative, and the two answers have imaginary numbers. But not this day.) You should have two real answers here.

OpenStudy (mjdennis):

(-4+ ___)/2 and (-4 - ___)/2 ??

OpenStudy (cherry1214):

idk im officially lost

OpenStudy (mjdennis):

Sorry. Ignore my long paragraph two comments back for now. Take my word for it, this problem will have two answers. Let's go back to the part where you gave me \[x= \frac{ -4 \pm \sqrt{4 ^{2} - 4(1)(3)} }{ 2(1) }\]

OpenStudy (mjdennis):

\[x= \frac{ -4 \pm \sqrt{4 ^{2} - 12} }{ 2 }\] One answer (call it x1, or x-sub-one), is: \[x_{1}= \frac{ -4 + \sqrt{4 ^{2} - 12} }{ 2 }\] The other is:\[x_{2}= \frac{ -4 - \sqrt{4 ^{2} - 12} }{ 2 }\]

OpenStudy (mjdennis):

My lunch break is over and I have to go. When you get your answers, try to put them back into the original equation x^2+3=-4x and see if both sides of the equation are equal.